OpenStudy (anonymous):

.5x + (-6) = 1 how do i find x???

OpenStudy (anonymous):

\[.5x -6 = 1~~~~ \implies .5x = 6+1\] \[.5x=7 \implies x = \frac{ 7 }{ .5 }\]

OpenStudy (anonymous):

you lost me Bats

OpenStudy (anonymous):

Alright lets draw it out then

OpenStudy (anonymous):

Easier to show all the cancellations and stuff

OpenStudy (anonymous):

i need to know what times one half + -6 euqals 1

OpenStudy (anonymous):

 With me so far?